A math professor notices that scores from a recent exam are normally distributed with a mean of 66 and a standard deviation of 4.
(a) What score do 25% of the students exam scores fall below? Enter your response to 2 decimal places.
Answer:
(b) Suppose the professor decides to grade on a curve. If the professor wants 0.15% of the students to get an A, what is the minimum score for an A? Enter your response rounded to the nearest INTEGER VALUE.
Answer:

Answer by stanbon(75887)

A math professor notices that scores from a recent exam are normally distributed with a mean of 66 and a standard deviation of 4.
(a) What score do 25% of the students exam scores fall below? Enter your response to 2 decimal places.
Answer: The z-score of that exam score is invNorm(0.25) = -0.6745
The exam score is -0.6745*4+ 66 = 63.30
-------------
(b) Suppose the professor decides to grade on a curve. If the professor wants
15% of the students to get an A, what is the minimum score for an A? Enter your response rounded to the nearest INTEGER VALUE.
Find the z-score of that student score:
z = invNorm(0.85) = 1.0364
Find the student score:: x = z*s+u = 1.0364*4+66 = 70 (when rounded)
-----
Cheers,
Stan H.
------------